OpenStudy (anonymous):

what multiplies to -168 and adds up to -22?

OpenStudy (anonymous):

xy = -168 x+y = -22 x=-y-22 (-y-22)(y)=-168 Solve the last for y and then plug it into either equation for x.

OpenStudy (anonymous):

what? ur going to in depth with the question! im looking for a simple awnser.

jimthompson5910 (jim_thompson5910):

use either trial and error (list out the factors of -168 and see which ones add to -22) or solve x^2 - 22x - 168 = 0 for x using the quadratic formula, the two results will help you find the two numbers.

OpenStudy (anonymous):

yeah ive been trying the trial and error method

jimthompson5910 (jim_thompson5910):

here are all the ways to multiply 2 whole numbers to get -168 1*(-168) = -168 2*(-84) = -168 3*(-56) = -168 4*(-42) = -168 6*(-28) = -168 7*(-24) = -168 8*(-21) = -168 12*(-14) = -168 (-1)*(168) = -168 (-2)*(84) = -168 (-3)*(56) = -168 (-4)*(42) = -168 (-6)*(28) = -168 (-7)*(24) = -168 (-8)*(21) = -168 (-12)*(14) = -168

OpenStudy (anonymous):

nice..

jimthompson5910 (jim_thompson5910):

1 + (-168) = -167 ... doesn't work since we want -22 2 + (-84) = -82 ... doesn't work since we want -22 Keep going until you find -22 as a result You may not find it (there's no guarantee it's in this list)

OpenStudy (anonymous):

I was trying to teach you MattEditz versus just telling you an answer so that you can ask the question very similar again in 10 minutes.

OpenStudy (anonymous):

well i already know how to do it, i just cant figure this one out.

OpenStudy (anonymous):

If you cannot figure it out you do not know how to do it. Accept their instruction and re-try for yourself.

OpenStudy (anonymous):

really? that makes no sense.

OpenStudy (anonymous):

It means that if you look at the current Open Questions half of them are the same as another question. And the other half have to do with interest.

OpenStudy (anonymous):

all u guys were useless

OpenStudy (anonymous):

He literally spelled out the answer for you, learn some respect.

OpenStudy (anonymous):

not really